# Isospin precession in

### v1

@@ I Introduction: the vector potential in gauge theory
electrically charged classical particles move freely outside a cylinder carrying a magnetic flux, however the curl-free vector potential $A=A_{\mu}dx^{\mu}$ allows for a quantum effect. The experimental confirmation of the Aharonov-Bohm prediction ABexp ; Osakabe ; Peshkin89 ; Tonomura87 ; Tonomura89 ; TonomuraToday led Wu and Yang to argue in their seminal paper WuYang75 that :
@@
Two vector potentials are therefore physically equivalent whenever their fluxes differ by an even multiple of $\pi$ AharonovBohm ; ABexp , $\Phi^{(1)}-\Phi^{(2)}={2N}\,\pi,\;N=0,\,\pm 1,\dots$ Peshkin89 ; Tonomura87 ; Tonomura89 ; TonomuraToday .
@@
In this paper we argue that while there is no classical electromagnetic Aharonov-Bohm effect, however in the non-Abelian  case a “shadow” of the quantum effect, namely isospin precession WuYang75 ; NABA82 ; MSW ; Jackiw86 ; KernerFest can be recovered at the (pseudo-) classical level. The difference with the Abelian case comes form that a particle with non-Abelian  structure carries an additional, internal degree of freedom represented by an isospin vector $\vec{Q}=(Q^{a})$ [or equivalently, by an ${\mathrm{su}}(2)$ matrix in the Lie algebra of the gauge group] which couples the particle to the YM field. The clue is that, unlike the electric charge, the isospin has non-trivial dynamics rather than being a mere constant Kerner68 ; Wong70 ; Trautman70 ; Cho75 ; Witten81 ; Balachandran76 ; Balachandran77 ; Sternberg77 ; Sternberg78 ; DuvalAix79 ; DH82 . Then our first result says:
@@
The original Wu-Yang proposal does not appear practical  [^6] Zeilinger et al. Zeilinger did scatter a neutron beam on a rotating rod of ${}^{138}$, setting up an upper limit $10^{-15}$ on the strength of the non-Abelian  effect when compared to its electromagnetic counterpart.. Therefore the Wu-Yang proposal has long be view as a Gedankenexperiment and the study was focused at particles coupled to a pure $\mathrm{SU}(2)$ Yang-Mills field with no electromagnetic interaction  [^7] Ignoring the electric charge is reminiscent of von Neumann’s “dry water”, though.. However, as emphasised by Yang and Mills already YangMills ,
@@
Then in sec.VI we derive conserved charges, associated with internal symmetries. Remarkably, non-Abelian  Aharonov-Bohm field admits two types of the latter: a trivial one and a “twisted” one, which arise when the enclosed flux is quantized in even resp. in odd integers, see eqn. (VI.1) HR-PRD ; HR-color ; NABAPRD .
Important conceptual and experimental progress has been made since those times, though WILCZEKCOLEMAN ; Wilczek89 ; Alford90 ; Preskill90 ; Brandenberger93 ; MSW ; Jackiw86 ; MoreNABA . The developments Osterloh ; Goldman ; Jacob ; Dalibard ; YChen ; YYang ; YBiao ; YangYang which concern effective non-Abelian gauge fields change the assessment dramatically, opening the way to laboratory testing.
This paper celebrates Tai-Tsun Wu’s 90th birthday. It was indeed “Tai” who, 42 years ago, suggested to one of us (PAH) to study the Non-Abelian Aharonov-Bohm Effect .
@@ II The Wu-Yang phase factor
$\vec{Q}\equiv{\tiny\left(\begin{array}[]{c}Q^{1}\\ Q^{2}\\ Q^{3}\end{array}\right)}\;\;\Leftrightarrow\;\;Q^{1}\,\frac{\sigma_{1}}{2i}+Q^{2}\,\frac{\sigma_{2}}{2i}+Q^{3}\,\frac{\sigma_{3}}{2i}=\hbox{$\textstyle\frac{1}{2i}$}{\tiny\left(\begin{array}[]{cc}Q^{3}&Q^{1}-iQ^{2}\\[6.0pt] Q^{1}+iQ^{2}&-Q^{3}\end{array}\right)}\equiv Q\,.$ (II.2)
@@
is a rotation by angle ${\Delta{\vartheta}}\,{\mathfrak{a}}$ around the 3rd axis in isospace. For a full circle we recover the Wu-Yang factor, $\mathcal{F}(2\pi)=\mathfrak{F}$ in (II). The non-Abelian Wu-Yang factor (II.6d) can be presented in terms of its Abelian counterparts (I.2),
@@
The additional concept we needed is that of classical isospin, represented by a 3-vector $\vec{Q}$ or alternatively, by an ${\mathrm{su}}(2)$ matrix, eqn. (II.2). Limiting ourselves at a non-relativistic study in flat space, the equations of motion of a classical isospin-carying particle put forward by Kerner, and by Wong, Kerner68 ; Wong70 and further discussed in Trautman70 ; Cho75 ; Witten81 ; Balachandran76 ; Balachandran77 ; Sternberg77 ; Sternberg78 ; DuvalAix79 ; DH82 are,
@@
where $Q_{0}$ is an arbitrary point in an adjoint orbit ${\mathcal{O}}=\big{\{}g^{-1}Q_{0}g\,|\,g\in G\big{\}}$ of the gauge group Sternberg77 ; Sternberg78 ; DuvalAix79 ; DH82 ; HPAAix79 . For $G=\mathrm{SU}(2)$ the orbit is a 2-sphere ${\mathds{S}}^{2}$ of radius $|Q_{0}|$ in isospace.
@@ III.1 Internal charges and symmetries
However diatomic molecules MSW ; Jackiw86 ; KernerFest provide us with an example where no covariantly constant direction fields and thus no conserved charges. Our investigations in sec.VI then show that this is also what happens in the Non-Abelian Aharonov-Bohm Effect — with the remarkable exception when the non-Abelian  flux is quantized as in (VI.1).
@@ III.2 Conserved quantities
Conserved quantities are usually derived from space-time symmetries by Noether’s theorem JackiwManton ; AbbottDeser . Below we argue that the charge (III.5) is generated by an internal symmetry HR-PRD ; HR-color . We first recall some general facts Schwarz77 ; Harnad79 ; ForgacsManton ; JackiwManton ; Jackiw80 ; DH82 . A vectorfield $X$ is a symmetry for a gauge field if the variation of the vector potential can be compensated by a gauge transformation,
@@
where $q$ is the conserved charge (III.8). This expression is analogous to the one in the field of a Dirac monopole JackiwRebbi76 ; Hasi ; JackiwManton ; Jackiw80 ; DH82 ; Jackiw86 ; vHolten ; HP-NGOME . Eliminating ${\vec{\pi}}$ in favor of $m\dot{{\vec{x}}}={\vec{p}}={\vec{\pi}}+{\vec{A}}$, (III.12) ${\vec{x}}\times{\vec{A}}$ and $q{\hat{\bm{r}}}$ combine into the isospin vector, and the angular momentum is separated into the sum of orbital and internal terms,
@@ IV Isospin precession
As emphasised in the introduction, there is no classical Aharonov-Bohm effect in the Abelian theory : charged particles move freely in the curl-free magnetic potential. The effect is purely quantum mechanical and is related to the phase freedom AharonovBohm ; ABexp ; Tonomura87 ; Peshkin89 ; Tonomura89 ; HPAAix79 .
@@
The classical dynamics is given by the Kerner-Wong equations (III). Outside the cylinder the Yang-Mills field is flat,  $F=DA=0$, therefore the generalized Lorentz force in (III.1a) vanishes and the space-time motion is free, just like in the Abelian case.
@@ V Isospin scattering
Eqns. (IV.6)-(IV.7) with $\Delta\theta=\pi$ allow us the express the result in terms of the Wu-Yang factor (II) in matrix resp. in vector form NABA82 ,
@@
$\mathfrak{F}=\left\{\begin{array}[]{rlc}\mathds{I}&{\quad\text{\small when}\quad}&{\mathfrak{a}}=2k\\ -\mathds{I}&{\quad\text{\small when}\quad}&{\mathfrak{a}}=1+2k\end{array}\right.\;.$ (V.4)
@@
The non-integrable phase factor for the half-loop (II.9) is the square root of the Wu-Yang factor (II),
@@
Figure 5: Isospin motion for ${\mathfrak{a}}=1$ for both protons (V) and neutrons (V), in both respective half-planes. The trajectories follow cycloid segments, which all satisfy ${\vec{Q}}_{out}=-{\vec{Q}}_{in}$.
@@
[…]
@@ VI.1 Implementability and internal symmetry
[…]
@@
[…]
@@
[…]
@@ VI.2 Conserved charges in the Non-Abelian Aharonov-Bohm Effect
${\widehat{\Psi}}_{\sigma_{3}}(\varphi)=$ $\;\;\quad\frac{1}{2i}{\tiny\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)}\quad{\quad\Rightarrow\quad}\;\;\,q_{vert}\;\;={Q^{3}(0)}\,.$ (VI.17f)
@@
$\left\{\begin{array}[]{llll}Q^{1}(\varphi)=\frac{\sigma_{1}}{2i}&&q^{1}=&\cos\varphi\neq\mathop{\rm const}\\[5.0pt] Q^{2}(\varphi)=\frac{\sigma_{2}}{2i}&&q^{2}=&\sin\varphi\neq\mathop{\rm const}\\[5.0pt] Q^{3}(\varphi)=\frac{\sigma_{3}}{2i}&&q^{3}=&0=\mathop{\rm const}\end{array}\right.$ (VI.29)
@@ VII Conserved quantities by van Holten’s algorithm
$\mu\to\left\{\begin{array}[]{lrlll}\mu_{+}=&\;{{\mathfrak{a}}\pi}/{2}+\phi_{0}&{\quad\text{\small when}\quad}&y>0\\[3.0pt] \mu_{-}=&-{{\mathfrak{a}}\pi}/{2}+\phi_{0}&{\quad\text{\small when}\quad}&y<0\end{array}\right.{\quad\Rightarrow\quad}\mu_{-}=\mu_{+}-{a\pi}\,.$ (VII.9)
@@
When the Wu-Yang factor does not belong to the center, $\mathfrak{F}\neq\pm\mathds{I}$, then only the 3rd component of the orbital angular momentum, $L=L_{3}$, is conserved. However when the flux is quantized in either of the two series in (VI.1), then $L_{3}$ and $Q^{3}$ in (IV.1) and (VI.32), are both separately conserved and the symmetry is doubled, (VI.33), consistently with Souriau’s décomposition barycentrique SSD .
@@ VIII Conclusion and outlook
This project, suggested by T. T. Wu to PAH over 40 years ago NABA82 ; NABAPRD ; HP-EPL ; HP-Kollar ; Sundrum 191919 H. Bacry, musing over a piece of paper in Marseille in the eighties, more than once whispered : “On écrit le même article durant toute sa vie” (One writes the same paper during all his life)., studies the Non-Abelian Aharonov-Bohm Effect proposed by Wu and Yang in their celebrated paper on the non-integrable phase factor WuYang75 . While the original proposal is field-theoretical/quantum mechanical, here we argue that a “shadow” of it, namely isospin precession [^20] Isospin precession was found also for diatomic molecules MSW ; Jackiw86 ; see the recent study KernerFest ., can be recovered also at the the pseudo-classical level, with the isospin represented by a vector ${\vec{Q}}$ of the Lie algebra of the gauge group  [^21] Subtleties of the concept of classical isospin were discussed in Arodz82 .. This is in sharp contrast with the Abelian (electromagnetic) case : Aharonov-Bohm effect is purely quantum mechanical AharonovBohm .
@@
In the non-Abelian  Aharonov-Bohm case, though, a covarantly constant ${\widehat{\Psi}}$ exists only when the non-Abelian  flux is quantized in either of the two distinct series (VI.1), then we get 3 conserved internal charges HR-PRD ; HR-color ; for other flux values we do not get any. The flux quantization we find here is reminiscent of the one in superconducting rings DeaverFairbank ; ByersYang ; Onsager .
@@
Such an eventuality raises the fundamental question : Is such a transition indeed possible or not ? What would happen with the electric charge ? What could be the physical mechanism for it. We confess not to be able to answer. It might well happen that the test particle framework we follow in this paper (or its quantum mechanical antecedents NABA82 ; HP-Kollar ; HP-EPL ) are just not sufficient to answer. Another idea could be to enlarge the commutative part of the gauge group from $\mathrm{SU}(2)$ to Weinberg-Salam $\mathrm{U}(2)$ (locally $(\mathrm{U}(2)\times\mathrm{U}(1)$).
@@
The angular momentum in the Non-Abelian Aharonov-Bohm Effect exhibits the “spin from isospin” phenomenon JackiwRebbi76 ; Hasi ; JackiwManton ; DH82 ; Chen:2008ag . The orbital angular momentum is conserved for all fluxes. However for the two particular values of the flux in (VI.1), we have additional terms generated by internal symmetry which is independent of space-time rotations. The total angular momentum, $J$ in (VI.30) splits into two separately conserved pieces, one external (orbital) and one internal (isospin), realizing Souriau’s doubled rotational symmetry idea SSD . When the flux quantized in odd units, eqn (VI.11b), the remnant of the internal symmetry is detected by the precessions of the isospin on horizontal circles, shown in FIG.6.
@@
Quantum aspects were studied in NABA82 ; HP-Kollar ; HR-PRD ; HR-color ; NABAPRD ; HP-EPL ; Sundrum . See MoreNABA for a (very incomplete) selection. It is remarkable that part of the Non-Abelian Aharonov-Bohm Effect can be captured by the (pseudo-)classical isospin of Kerner and Wong Kerner68 ; Wong70 , confirming that the isospin is neither a truly classical nor a truly quantum concept: it is at half ways between the two, justifying our terminology.
Yang et al YYang ; MITNews attribute the non-Abelian  Aharonov-Bohm effect to the breaking of time-reversal symmetry. However parallel transport implies that real time is replaced, in our framework, by the angle $\varphi$. The effect thus comes from “navigating” on either sides of the flux. However we found that when the flux is quantized, (VI.1) then the time reversal symmetry disappears: side-dependent phase shifts yield identical contributions on both sides, as highlighted by doubling the rotational symmetry, as in (VI.30)-(VI.33).
Ongoing work on effective/artifical gauge fields MSW ; Jackiw86 ; KernerFest ; MoreNABA ; Osterloh ; Goldman ; Jacob ; Dalibard ; YChen ; YYang ; YBiao ; YangYang provide promising testing ground of the conceptual suggestions of Wu and Yang, allowing us to hope that the Non-Abelian Aharonov-Bohm Effect may, on one day, not be considered anymore as …“dry water”.

### v3

@@ I Introduction: the vector potential in gauge theory
electrically charged classical particles move freely outside a cylinder carrying a magnetic flux, however the curl-free vector potential $A=A_{\mu}dx^{\mu}$ allows for a quantum effect. The experimental confirmation of the Aharonov-Bohm prediction Chambers ; Mollen ; Osakabe ; Peshkin89 ; Tonomura87 ; Tonomura89 ; TonomuraToday led Wu and Yang to argue in their seminal paper WuYang75 that :
@@
Two vector potentials are therefore physically equivalent whenever their fluxes differ by an even multiple of $\pi$ AharonovBohm ; Chambers ; Mollen , $\Phi^{(1)}-\Phi^{(2)}={2N}\,\pi,\;N=0,\,\pm 1,\dots$ Peshkin89 ; Tonomura87 ; Tonomura89 ; TonomuraToday .
@@
In this paper we argue that while there is no classical electromagnetic Aharonov-Bohm effect, however in the non-Abelian  case a “shadow” of the quantum effect, namely isospin precession WuYang75 ; NABA82 ; MSW ; Jackiw86 ; KernerFest can be recovered at the (pseudo-) classical level. The difference with the Abelian case comes from that a particle with non-Abelian  structure carries an additional, internal degree of freedom represented by an isospin vector $\vec{Q}=(Q^{a})$ [or equivalently, by a matrix in the Lie algebra, ${\mathrm{su}}(2)$, of the gauge group] which couples the particle to the YM field. The clue is that, unlike the electric charge, the isospin has non-trivial dynamics rather than being a mere constant Kerner68 ; Wong70 ; Trautman70 ; Cho75 ; Witten81 ; Balachandran76 ; Balachandran77 ; Sternberg77 ; Sternberg78 ; DuvalAix79 ; DH82 . Then our first result says:
@@
The original Wu-Yang proposal does not appear practical  [^6] Zeilinger et al. Zeilinger did actually scatter a neutron beam on a rotating rod of 138U, setting up an upper limit $10^{-15}$ on the strength of the non-Abelian  effect when compared to its electromagnetic counterpart.. Therefore the Wu-Yang proposal has long be view as a Gedankenexperiment and the study was focused at particles coupled to a pure $\mathrm{SU}(2)$ Yang-Mills field with no electromagnetic interaction  [^7] Ignoring the electric charge is reminiscent of von Neumann’s “dry water”, though.. However, as emphasised by Yang and Mills already YangMills ,
@@
Then in sec.VI we derive conserved charges, associated with internal symmetries. Remarkably, non-Abelian  Aharonov-Bohm field admits two types of the latter: a trivial one and a “twisted” one, which arise when the enclosed flux is quantized in even resp. in odd integers, see eqn. (VI.11) HR-PRD ; HR-color ; NABAPRD .
Important conceptual and experimental progress has been made since those times, though WILCZEKCOLEMAN ; Wilczek89 ; Alford90 ; Preskill90 ; Brandenberger93 ; MSW ; Jackiw86 ; KernerFest ; Bright15 ; Hosseini16 ; Chatterjee15 ; Cserti . The developments Fruchart ; Osterloh ; Goldman ; Jacob ; Dalibard ; YChen ; YYang ; YBiao ; YangYang ; NABA-PR which concern effective non-Abelian gauge fields change the assessment dramatically, opening the way to laboratory testing.
@@ II The Wu-Yang phase factor
$\vec{Q}\equiv{\tiny\left(\begin{array}[]{c}Q^{1}\\ Q^{2}\\ Q^{3}\end{array}\right)}\;\;\Leftrightarrow\;\;Q^{1}\,\frac{\sigma_{1}}{2i}+Q^{2}\,\frac{\sigma_{2}}{2i}+Q^{3}\,\frac{\sigma_{3}}{2i}=\hbox{$\textstyle\frac{1}{2i}$}{\left(\begin{array}[]{cc}Q^{3}&Q^{1}-iQ^{2}\\[6.0pt] Q^{1}+iQ^{2}&-Q^{3}\end{array}\right)}\equiv Q\,.$ (II.2)
@@
is a rotation by angle ${\Delta{\vartheta}}\,{\mathfrak{a}}$ around the 3rd axis in isospace. For a full circle we recover the Wu-Yang factor, $\mathcal{F}(2\pi)=\mathfrak{F}$ in (II.6). The non-Abelian Wu-Yang factor (II.6d) can be presented in terms of its Abelian counterparts (I.2),
@@
The additional concept we needed is that of classical isospin, represented by a 3-vector $\vec{Q}$ or alternatively, by an ${\mathrm{su}}(2)$ matrix, eqn. (II.2). Limiting ourselves at a non-relativistic study in flat space, the equations of motion of a classical isospin-carying particle put forward by Kerner Kerner68 , and by Wong Wong70 , and further discussed in Trautman70 ; Cho75 ; Witten81 ; Balachandran76 ; Balachandran77 ; Sternberg77 ; Sternberg78 ; DuvalAix79 ; Duval78 ; DH82 ; Wipf85 ; Feher86 are,
@@
where $Q_{0}$ is an arbitrary point in an adjoint orbit ${\mathcal{O}}=\big{\{}g^{-1}Q_{0}g\,|\,g\in G\big{\}}$ of the gauge group Sternberg77 ; Sternberg78 ; DuvalAix79 ; Duval78 ; DH82 ; Wipf85 ; Feher86 ; HPAAix79 . For $G=\mathrm{SU}(2)$ the orbit is a 2-sphere ${\mathds{S}}^{2}$ of radius $|Q_{0}|$ in isospace.
@@ III.1 Internal charges and symmetries
However diatomic molecules MSW ; Jackiw86 ; KernerFest provide us with an example where no covariantly constant direction fields and thus no conserved charges. Our investigations in sec.VI then show that this is also what happens in the Non-Abelian Aharonov-Bohm Effect — with the remarkable exception when the non-Abelian  flux is quantized as in (VI.11).
@@ III.2 Conserved quantities
Conserved quantities are usually derived from space-time symmetries by Noether’s theorem JackiwManton ; AbbottDeser . Below we argue that the charge (III.5) is generated by an internal symmetry HR-PRD ; HR-color . We first recall some general facts Schwarz77 ; Harnad79 ; ForgacsManton ; JackiwManton ; Jackiw80 ; DH82 ; Wipf85 ; Feher86 . A vectorfield $X$ is a symmetry for a gauge field if the variation of the vector potential can be compensated by a gauge transformation,
@@
where $q$ is the conserved charge (III.8). This expression is analogous to the one in the field of a Dirac monopole JackiwRebbi76 ; Hasi ; JackiwManton ; Jackiw80 ; DH82 ; Wipf85 ; Feher86 ; Jackiw86 ; vHolten ; HP-NGOME . Eliminating ${\vec{\pi}}$ in favor of $m\dot{{\vec{x}}}={\vec{p}}={\vec{\pi}}+{\vec{A}}$, (III.12) ${\vec{x}}\times{\vec{A}}$ and $q{\hat{\bm{r}}}$ combine into the isospin vector, and the angular momentum is separated into the sum of orbital and internal terms,
@@ IV Isospin precession
As emphasised in the introduction, there is no classical Aharonov-Bohm effect in the Abelian theory : charged particles move freely in the curl-free magnetic potential. The effect is purely quantum mechanical and is related to the phase freedom AharonovBohm ; Chambers ; Mollen ; Tonomura87 ; Peshkin89 ; Tonomura89 ; HPAAix79 .
@@
The classical dynamics is given by the Kerner-Wong equations (III.1). Outside the cylinder the Yang-Mills field is flat,  $F=DA=0$, therefore the generalized Lorentz force in (III.1a) vanishes and the space-time motion is free, just like in the Abelian case.
@@ V Isospin scattering
Eqns. (IV.6)-(IV.7) with $\Delta\theta=\pi$ allow us the express the result in terms of the Wu-Yang factor (II.6) in matrix resp. in vector form NABA82 ,
@@
$\mathfrak{F}=\left\{\begin{array}[]{rlc}\mathds{I}&{\quad\text{\small when}}&{\mathfrak{a}}=2k\\ -\mathds{I}&{\quad\text{\small when}}&{\mathfrak{a}}=1+2k\end{array}\right.\;.$ (V.4)
@@
The non-integrable phase factor for the half-loop (II.9) is the square root of the Wu-Yang factor (II.6),
@@
Figure 5: Isospin motion for ${\mathfrak{a}}=1$ for both protons (V.6) and neutrons (V.7), in both respective half-planes. The trajectories follow cycloid segments, which all satisfy ${\vec{Q}}_{out}=-{\vec{Q}}_{in}$.
@@
$\vec{Q}_{out}=S\,\vec{Q}_{in}{\quad\text{\small where}\quad}S=R_{\pm{\pi}{\mathfrak{a}}},$ (V.10)
@@ VI.1 Implementability and internal symmetry
$\Omega_{k}^{(1)}(x)=k=\big{(}k_{ij}\big{)}\;{\quad\text{\small and}\quad}\,\;\Omega_{k}^{(2)}(x)=\left(\begin{array}[]{cc}k_{11}&e^{i\varphi}k_{12}\\ e^{-i\varphi}\bar{k}_{12}&\bar{k}_{11}\end{array}\right)\,,$ (VI.3)
@@
${\widehat{\Psi}}_{\kappa}^{(1)}(\varphi)=\;{\frac{1}{2i}}\kappa={\frac{1}{2i}}\big{(}\kappa_{ab}\big{)}\,{\quad\text{\small and}\quad}\,{\widehat{\Psi}}_{\kappa}^{(2)}(\varphi)={\frac{1}{2i}}\left(\begin{array}[]{cc}\kappa_{11}&e^{i\varphi}\,\kappa_{12}\\[2.0pt] e^{-i\varphi}\,\overline{\kappa}_{12}&-\kappa_{11}\end{array}\right)\,,$ (VI.4)
@@
$\begin{array}[]{lclc}q_{b}^{(1)}=&\quad Q^{b}&{\quad\text{\small for}}&{\mathfrak{a}}=2k\\[6.0pt] q_{b}^{(2)}=&\;-2\mathrm{tr}\left(Q\cdot{{\widehat{\Psi}}^{(2)}}_{\sigma_{b}}\right)&{\quad\text{\small for}}&{\mathfrak{a}}=1+2k\end{array}$ (VI.13)
@@ VI.2 Conserved charges in the non-Abelian Aharonov-Bohm effect
${\widehat{\Psi}}_{\sigma_{3}}(\varphi)=$ $\;\;\quad\frac{1}{2i}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\quad{\quad\Rightarrow\quad}\;\;\,q_{vert}\;\;={Q^{3}(0)}\,.$ (VI.17f)
@@
$\left\{\begin{array}[]{llll}Q^{1}(\varphi)=\frac{\sigma_{1}}{2i}&&q^{1}=&\cos\varphi\neq\mathop{\rm const}\nolimits\\[5.0pt] Q^{2}(\varphi)=\frac{\sigma_{2}}{2i}&&q^{2}=&\sin\varphi\neq\mathop{\rm const}\nolimits\\[5.0pt] Q^{3}(\varphi)=\frac{\sigma_{3}}{2i}&&q^{3}=&0=\mathop{\rm const}\nolimits\end{array}\right.$ (VI.29)
@@ VII Conserved quantities by van Holten’s algorithm
$\mu\to\left\{\begin{array}[]{lrlll}\mu_{+}=&\;{{\mathfrak{a}}\pi}/{2}+\phi_{0}&{\quad\text{\small when}}&y>0\\[3.0pt] \mu_{-}=&-{{\mathfrak{a}}\pi}/{2}+\phi_{0}&{\quad\text{\small when}}&y<0\end{array}\right.{\quad\Rightarrow\quad}\mu_{-}=\mu_{+}-{a\pi}\,.$ (VII.9)
@@
When the Wu-Yang factor does not belong to the center, $\mathfrak{F}\neq\pm\mathds{I}$, then only the 3rd component of the orbital angular momentum, $L=L_{3}$, is conserved. However when the flux is quantized in either of the two series in (VI.11), then $L_{3}$ and $Q^{3}$ in (IV.1) and (VI.32), are both separately conserved and the symmetry is doubled, (VI.33), consistently with Souriau’s décomposition barycentrique SSD .
@@ VIII Conclusion and outlook
This project, suggested by T. T. Wu to PAH over 40 years ago NABA82 ; NABAPRD ; HP-EPL ; HP-Kollar1 ; HP-Kollar2 ; Sundrum 191919 H. Bacry, musing over a piece of paper in Marseille in the eighties, more than once whispered : “On écrit le même article durant toute sa vie” (One writes the same paper during all his life)., studies the Non-Abelian Aharonov-Bohm Effect proposed by Wu and Yang in their celebrated paper on the non-integrable phase factor WuYang75 . While the original proposal is field-theoretical/quantum mechanical, here we argue that a “shadow” of it, namely isospin precession [^20] Isospin precession was found also for diatomic molecules MSW ; Jackiw86 ; see the recent study KernerFest ., can be recovered also at the the pseudo-classical level, with the isospin represented by a vector ${\vec{Q}}$ of the Lie algebra of the gauge group  [^21] Subtleties of the concept of classical isospin were discussed in Arodz82 ; Arodz82b ; Arodz83 .. This is in sharp contrast with the Abelian (electromagnetic) case : Aharonov-Bohm effect is purely quantum mechanical AharonovBohm .
@@
In the non-Abelian  Aharonov-Bohm case, though, a covarantly constant ${\widehat{\Psi}}$ exists only when the non-Abelian  flux is quantized in either of the two distinct series (VI.11), then we get 3 conserved internal charges HR-PRD ; HR-color ; for other flux values we do not get any. The flux quantization we find here is reminiscent of the one in superconducting rings DeaverFairbank ; ByersYang ; Onsager .
@@
Such an eventuality raises the fundamental question : Is such a transition indeed possible or not ? What would happen with the electric charge ? What could be the physical mechanism for it. We confess not to be able to answer. It might well happen that the test particle framework we follow in this paper (or its quantum mechanical antecedents NABA82 ; HP-Kollar1 ; HP-Kollar2 ; HP-EPL ) are just not sufficient to answer. Another idea could be to enlarge the commutative part of the gauge group from $\mathrm{SU}(2)$ to Weinberg-Salam $\mathrm{U}(2)$ (locally $(\mathrm{U}(2)\times\mathrm{U}(1)$).
@@
The angular momentum in the Non-Abelian Aharonov-Bohm Effect exhibits the “spin from isospin” phenomenon JackiwRebbi76 ; Hasi ; JackiwManton ; DH82 ; Chen:2008ag . The orbital angular momentum is conserved for all fluxes. However for the two particular values of the flux in (VI.11), we have additional terms generated by internal symmetry which is independent of space-time rotations. The total angular momentum, $J$ in (VI.30) splits into two separately conserved pieces, one external (orbital) and one internal (isospin), realizing Souriau’s doubled rotational symmetry idea SSD . When the flux quantized in odd units, eqn (VI.11b), the remnant of the internal symmetry is detected by the precessions of the isospin on horizontal circles, shown in FIG.6.
@@
Quantum aspects were studied in NABA82 ; HP-Kollar1 ; HP-Kollar2 ; HR-PRD ; HR-color ; NABAPRD ; HP-EPL ; Sundrum . See KernerFest ; Bright15 ; Hosseini16 ; Chatterjee15 ; Cserti for a (very incomplete) selection. It is remarkable that part of the Non-Abelian Aharonov-Bohm Effect can be captured by the (pseudo-)classical isospin of Kerner and Wong Kerner68 ; Wong70 , confirming that the isospin is neither a truly classical nor a truly quantum concept: it is at half ways between the two, justifying our terminology.
Yang et al YYang ; MITNews attribute the non-Abelian  Aharonov-Bohm effect to the breaking of time-reversal symmetry. However parallel transport implies that real time is replaced, in our framework, by the angle $\varphi$. Thus the effect comes from “navigating” on either sides of the flux. However we found that when the flux is quantized, (VI.11) then the time reversal symmetry disappears: side-dependent phase shifts yield identical contributions on both sides, as highlighted by doubling the rotational symmetry, as in (VI.30)-(VI.33).
Intensive ongoing work on effective/artifical gauge fields MSW ; Jackiw86 ; KernerFest ; Bright15 ; Hosseini16 ; Chatterjee15 ; Cserti ; Osterloh ; Goldman ; Jacob ; Dalibard ; YChen ; YYang ; Fruchart ; YBiao ; YangYang ; NABA-PR provide promising testing ground of the conceptual suggestions of Wu and Yang, allowing us to hope that the Non-Abelian Aharonov-Bohm Effect may, on one day, not be considered anymore as …“dry water” (using a celebrated expression put forward by J. von Neumann).